OpenStudy (anonymous):

Jack usually mows his lawn in 7 hours. Marilyn can mow the same yard in 4 hours. How much would it take for them to mow the lawn together?

OpenStudy (anonymous):

Jack: Time = 7 hrs/job ; rate = 1/7 job/hr Marilyn: Marilyn = 4 hrs/job ; rate = 1/4 job/hr Together: Time = x hrs/job ; rate = 1/x job/hr EQUATION: rate + rate = together rate 1/7 + 1/4 = 1/x 11/28 = 1/x x = 28/11 = 2.545454... hrs